# Cohet: A CXL-Driven Coherent Heterogeneous Computing Framework with Hardware-Calibrated Full-System Simulation

### v1

@@ V-A Remote Atomic Operations
Atomic operations play a pivotal role in parallel programming, as they are widely used to ensure consistency of shared data accesses [dynamo] and implement lock-free algorithms [non_lock] as well as synchronization primitives such as spinlocks [atomic_cache, lock_1, lock_2]. As a result, modern processors have been highly optimized for efficient atomic operations within a single multi-core node. However, the emergence of data-intensive workloads (e.g., machine learning and graph analytics) has driven the distribution of massive shared data sets across multiple computing nodes in order to exploit data parallel processing. This shift gives rise to the demand for frequent cross-node remote atomic operations (RAOs)[circustent, RAE, rao_sequencer, think_more_rdma, rdma_design, lock_2, ao_evaluating]. RAOs enforce atomic updates to shared data residing in a remote memory location, which is an extension of local atomics to distributed computing systems by means of remote direct memory access (RDMA). The study in [RAE] shows that when the number of computing nodes exceeds 64 and global memory accesses follow a random indexing pattern, RAOs can constitute as much as 98.44% of all atomic operations. Hence, the design and implementation efficiency of RAOs are crucial to the performance of large-scale high-performance computing systems.
@@ V-A1 Conventional RAO offloading design in PCIe-NIC
Traditional software-based implementations of RAOs rely on parallel programming libraries such as MPI and OpenSHMEM, but these implementations often suffer from performance overhead due to redundant software routines (e.g., multi-layer communication protocol stacks, software locks/semaphores, and redundant memory copies) [circustent, RAE]. To address this problem, modern supercomputers typically employ hardware offloading techniques in RDMA networks [infiniband] to optimize common RAO primitives such as compare-and-swap (CAS) and fetch-and-add (FAA). In such systems, RAOs are offloaded to the RDMA network interface controllers (NICs) and executed as indivisible read-modify-write operations across the PCIe interface, significantly reducing inter-node traffics. However, PCIe-NICs still face critical limitations when accelerating RAOs. That is, PCIe does not intrinsically guarantee memory consistency and atomicity, as well as timely visibility of data modifications to the host. Each RAO requires two DMA transfers, one for the read and another for the write, which must be issued consecutively without interruption for the same target address. This can be explained with an example illustrated in Fig. 8(a). When three FAA requests on a shared int variable are queued for processing, the NIC must launch six costly DMA transactions across PCIe. Moreover, due to PCIe’s relaxed ordering and split-transaction mechanisms [introduction_cxl], a later read request may arrive before a prior write, leading to potential read-after-write (RAW) hazards. To avoid such hazards, each RAO must wait for an acknowledgment of the previous write before proceeding, which significantly limits the RAO throughput [free_atomic].
@@ V-A2 Cohet-driven RAO offloading design in CXL-NIC
In contrast, Fig. 8(b) shows the CXL-based RAO design guided by Cohet to address this bottleneck. The CXL-NIC and the host CPU share a unified memory view. This allows the CXL-NIC to cache the target variable in the HMC, enabling subsequent RAOs to be serviced directly within the HMC without redundant PCIe transfers. The processing element (PE) locks the target RAO cacheline to prevent any invalidation during the RMW operation and thus preserve atomicity. Meanwhile, hardware-maintained cache coherence ensures that the host always perceives and retrieves the most recent data. This mechanism substantially improves the efficiency of RAO-intensive operations such as sequencers [rao_sequencer], lock services [lock_1, lock_2], and synchronization barriers [atomic_cache], which often involve many-to-one contention on a small set of hot memory locations. Furthermore, most data-intensive applications use pointer-based data structures such as graphs, unbalanced trees, unstructured grids, and sparse matrices. These structures generate fine-grained and irregular memory access patterns [graphpim]. As discussed in Section II-A, traditional DMA mechanisms are unsuitable for such access patterns. In contrast, CXL.cache efficiently supports such workloads through low-latency, coherent access to shared memory.
@@ V-B Remote Procedure Calls
Remote procedure calls (RPCs) serve as the fundamental inter-process communication layer for modern cloud services, supporting diverse distributed applications (e.g., microservices [micro_server], cloud storage [cloud_storage], and machine learning [ML]). An RPC abstracts remote compute resources by making a function call to a remote machine similar to a local call. Its underlying stack (typically implemented as a user-space library [grpc, thrift]) transparently handles connection management, network protocols, parameter (de)serialization, and thread scheduling. Industry measurements reveal significant RPC overheads: Google reports that RPCs account for 7.1% of CPU cycles in production clusters [google_2023]; Meta observes that parameter (de)serialization alone consumes 6.7% of CPU cycles in seven critical microservices [meta_2020]. These findings highlight the importance of accelerating the RPC stack especially (de)serialization to reclaim CPU resources for application workloads.
@@ V-B1 Conventional RPC offloading design in PCIe-NIC
Among existing serialization libraries [noauthor_protocol_nodate, noauthor_capn_nodate, noauthor_flatbuffers_nodate], we focus on Google’s Protobuf [noauthor_protocol_nodate] in this paper, as it has widespread adoption in today’s cloud applications. Two key characteristics of RPC messages critically impact acceleration designs. (1) The vast majority of messages are extremely small; a Google cluster analysis shows that 56% of messages are bytes and 93% are bytes [protoAcc]. (2) Nested messages can be defined via pointer references, analogous to pointer chasing, incurring significant cumulative overhead during (de)serialization[micro_server]; in real-world applications, the nesting of RPC message may exceed ten levels [protoAcc, google_2023].
A recent study RpcNIC [rpcnic] proposed a hardware offloading design of the RPC stack to a PCIe-NIC, which we have implemented in SimCXL as shown in Fig. 10. The host pre-runs the Protobuf compiler to store message structure metadata in a schema table, which guides message fields to decode in in-memory C++ objects or encode them into binary sequences. When an RPC request arrives, the NIC deserializer decodes it field-by-field, accumulating results in a on-chip temp buffer. When the current request’s deserialization is finished or the buffer is full, a one-shot DMA transfer to host memory is triggered. The NIC then increments the head pointer of a ring buffer in main memory through a DMA write to signal the arrival of a new request; the CPU processes the request and increments the tail pointer to indicate completion. For RPC responses, RpcNIC employs a pre-serialization mechanism; the CPU first invokes the on-chip memory copy engine (i.e., Intel’s DSA) to iteratively copy noncontiguous fields into a pre-allocated DMA-safe buffer. The CPU then updates an NIC-resident ring buffer via MMIO to signal the completion of pre-serialization. The NIC copies the prepared data from the host memory via a DMA read. The hardware serializer iteratively encodes the data in the NIC memory and transmits the RPC response over the network.
@@ VI-A1 Hardware testbed
Our hardware testbed is a high-performance dual-socket server, detailed in Table I. Each socket houses an Intel Gen4 Xeon Platinum 8468V processor, which integrates four CPU chiplets with each paired with a memory controller driving two memory channels. Each channel is populated with two DDR5 DIMMs. The four chiplets can be configured as a monolithic CPU or two/four NUMA nodes in sub-NUMA cluster mode (SNC) [4th_Xeon]. To ensure accurate and stable D2H memory access measurements, we enabled SNC-4 in the BIOS, dividing each socket into four separate NUMA nodes. Memory accesses from the same physical core to different NUMA nodes incur varying latencies depending on the number of NoC and UPI routing hops, thus manifesting the NUMA effect. The CXL devices on the testbed include an Intel Agilex I-Series FPGA Development Kit [agilex] and a Samsung memory expander [samsung_cmm_d]. The FPGA operates at and supports both CXL type-1/2 device configurations (denoted as CXL-FPGA) and standard PCIe configurations (denoted as PCIe-FPGA), connected to Socket1 via PCIe5.0 ×16. The CXL-FPGA integrates a hard CXL IP [cxl_fpga_ip] and is equipped with a 4-way HMC, whereas the PCIe-FPGA integrates a multi-channel DMA IP [multichannel_dma]. The Samsung expander, connected to Socket1 via PCIe5.0 ×8, is exposed to the system as a CPU-less NUMA node.
@@ VI-A2 SimCXL simulator
SimCXL is developed based on gem5 v23.1, with X86O3CPU and of DDR5 dual-channel memory. In terms of CXL devices, the device driver is developed on top of the Linux v6.12 kernel. The modular CXL controller in SimCXL can model CXL type-1/2/3 devices. Device memory can directly leverage gem5’s native DDR, NVM, and HBM models; device compute units are customized following the gem5 programming framework. For type-1/2 devices (denoted as CXL-FPGAsim), a 4-way HMC is implemented in the Ruby subsystem. Based on the latency breakdowns presented in the CXL specification [CXL_spec] and related studies [introduction_cxl, demystifying_type2, demystifying_type3, exploring_cxl_asic], the CXL controller and Ruby’s heterogeneous MESI protocol set multiple configurable parameters for tuning H2D and D2H access latency and bandwidth. For PCIe devices (denoted as PCIe-FPGAsim), we implemented DMA read and write engines to handle H2D and D2H data transfers, respectively. Both the CXL and PCIe device models can be configured to to match the FPGA configurations. Meanwhile, the frequency can also be set to to model real-world ASIC devices (denoted as CXL-ASICsim and PCIe-ASICsim, respectively), by frequency scaling based on the clock cycles measured from CXL-FPGA tests; this allows us to evaluate the realistic offloading performance of RAO and RPC on future production-grade CXL devices.
@@ VI-A4 Hardware calibration methodology
In this work, we calibrate the latency and bandwidth of CXL.cache for D2H memory accesses. The calibration of H2D memory access via CXL.mem will be presented in another paper. By tuning the model parameters of SimCXL, we are able to calibrate it to make its performance match that of CXL-FPGA. Note that the DMA performance calibration for SimCXL and PCIe-FPGA follows a similar approach. For D2H accesses via CXL.cache, cachelines may hit in the device’s HMC, host LLC, or host memory. HMC hits are tested by repeating address sequences. For LLC hits and memory hits, after initializing the target host memory addresses, we use the CLDEMOTE [cldemote] instruction to push cachelines to the LLC or the CLFLUSH [clflush] instruction to flush cachelines to memory, followed by the LSU test sequence. For predictable performance, hyper-threading and hardware prefetchers are disabled, and CPU frequency is fixed at 2.4GHz during test.
@@ VI-C SimCXL Bandwidth Calibration
To measure the bandwidth of CXL.cache, we configured the LSU to issue a large number of memory requests until a stable bandwidth value was achieved. We observed that when the request count exceeds 512 (totaling ), the bandwidths for HMC hit, LLC hit, and memory hit all converge. We then report the bandwidth results in Fig. 15 at 2,048 requests (totaling ), which are averaged over 1,000 tests. On the 400 MHz CXL-FPGA, the peak bandwidths for HMC hit, LLC hit, and memory hit stabilize at , , and , respectively. Given the FPGA’s theoretical maximum bandwidth of at , LLC and memory hits reach only 55% and 52.7% of the peak bandwidth. This degradation probably stems from pipeline bubbles introduced by cache-coherence checks during requests routed to the host, a behavior also observed in CPU accesses to remote NUMA nodes [demystifying_type3, exploring_cxl_asic]. In contrast, HMC accesses, eliminating the need for host-side coherence checks, achieve 97.7% of the theoretical bandwidth. One can see that SimCXL accurately reproduces the bandwidth of CXL-FPGA at .
@@ VI-D Evaluation of Remote Atomic Operations
We evaluated six RAO access patterns from the CircusTent benchmark [circustent] on both PCIe-NIC and CXL-NIC in SimCXL, with throughput speedups shown in Fig. 17. The CENTRAL pattern, a many-to-one workload that models a distributed lock service, achieves the highest speedup (40.2) by caching the hotspot data in CXL-NIC’s HMC, avoiding costly PCIe DMA transfers. STRIDE1 follows with a 22.4 speedup, as its sequential atomic updates on elements leverage cacheline granularity, enabling cache hits for seven subsequent operations after fetching a line. SCATTER and GATHER patterns, which involve randomized updates in the global address space, exhibit moderate speedups due to lower cache hit rates. The SG workload, which combines SCATTER and GATHER, follows a similar trend. Even the fully random pattern RAND, with a near-zero cache hit rate, delivers a 5.5 throughput gain due to the lower access latency of CXL.cache compared to PCIe-based DMA for fine-grained accesses. These results demonstrate that CXL-NIC significantly enhances the distributed RAO throughput across various access patterns by leveraging low-latency memory access and on-device caching of hot data.
@@ VI-E Evaluation of Remote Procedure Calls
We evaluated de/serialization time for CXL-based RPC (CXL-NIC.*) versus PCIe-based RPC (RpcNIC [rpcnic]) using HyperProtoBench [protoAcc]. Fig. 18(a) shows deserialization results, with CXL-NIC achieving speedups of 1.33 (Bench5) to 2.05 (Bench1) in six benchmarks. This acceleration comes primarily from the CXL-NIC’s low-latency fine-grained writes. Although Bench5’s large string fields favor DMA, CXL.cache still outperforms RpcNIC by eliminating redundant copies and software overhead. Bench1, with small-field messages, achieves the highest speedup. Moreover, CXL-NIC uses NC-P instructions to push fields directly into the LLC, improving subsequent CPU access performance.
@@
| Simulator/Emulator |
Full System |
Configurability | Sim. Error | Sim. Speed | ||||||||||
| CXLMemSim [CXLMemSim_github] | No | No | No | No | No | No | Medium | High | Medium | ||||||||||
| CXL-DMSim [cxl_dmsim] | No | No | Yes | No | Yes | Yes | High | Low | Low | ||||||||||
| Mess+gem5 [mess] | No | No | No | No | No | No | High | Medium | Low | ||||||||||
| QEMU [QEMU] | No | No | Yes | No | Yes | No | High | High | High | ||||||||||
| Remote NUMA [tarot_row_hammer, dmrpc] | No | No | No | No | No | N/A | Low | High | High | ||||||||||

### v2

@@ V-A Remote Atomic Operations
Atomic operations play a pivotal role in parallel programming, as they are widely used to ensure consistency of shared data accesses [75] and implement lock-free algorithms [50] as well as synchronization primitives such as spinlocks [99, 101, 56]. As a result, modern processors have been highly optimized for efficient atomic operations within a single multi-core node. However, the emergence of data-intensive workloads (e.g., machine learning and graph analytics) has driven the distribution of massive shared data sets across multiple computing nodes in order to exploit data parallel processing. This shift gives rise to the demand for frequent cross-node remote atomic operations (RAOs)[88, 85, 11, 49, 38, 56, 68]. RAOs enforce atomic updates to shared data residing in a remote memory location, which is an extension of local atomics to distributed computing systems by means of remote direct memory access (RDMA). The study in [85] shows that when the number of computing nodes exceeds 64 and global memory accesses follow a random indexing pattern, RAOs can constitute as much as 98.44% of all atomic operations. Hence, the design and implementation efficiency of RAOs are crucial to the performance of large-scale high-performance computing systems.
@@ V-A1 Conventional RAO offloading design in PCIe-NIC
Traditional software-based implementations of RAOs rely on parallel programming libraries such as MPI and OpenSHMEM, but these implementations often suffer from performance overhead due to redundant software routines (e.g., multi-layer communication protocol stacks, software locks/semaphores, and redundant memory copies) [88, 85]. To address this problem, modern supercomputers typically employ hardware offloading techniques in RDMA networks [35] to optimize common RAO primitives such as compare-and-swap (CAS) and fetch-and-add (FAA). In such systems, RAOs are offloaded to the RDMA network interface controllers (NICs) and executed as indivisible read-modify-write operations across the PCIe interface, significantly reducing inter-node traffics. However, PCIe-NICs still face critical limitations when accelerating RAOs. That is, PCIe does not intrinsically guarantee memory consistency and atomicity, as well as timely visibility of data modifications to the host. Each RAO requires two DMA transfers, one for the read and another for the write, which must be issued consecutively without interruption for the same target address. This can be explained with an example illustrated in Fig. 8(a). When three FAA requests on a shared int variable are queued for processing, the NIC must launch six costly DMA transactions across PCIe. Moreover, due to PCIe’s relaxed ordering and split-transaction mechanisms [70], a later read request may arrive before a prior write, leading to potential read-after-write (RAW) hazards. To avoid such hazards, each RAO must wait for an acknowledgment of the previous write before proceeding, which significantly limits the RAO throughput [5].
@@ V-A2 Cohet-driven RAO offloading design in CXL-NIC
In contrast, Fig. 8(b) shows the CXL-based RAO design guided by Cohet to address this bottleneck. The CXL-NIC and the host CPU share a unified memory view. This allows the CXL-NIC to cache the target variable in the HMC, enabling subsequent RAOs to be serviced directly within the HMC without redundant PCIe transfers. The processing element (PE) locks the target RAO cacheline to prevent any invalidation during the RMW operation and thus preserve atomicity. Meanwhile, hardware-maintained cache coherence ensures that the host always perceives and retrieves the most recent data. This mechanism substantially improves the efficiency of RAO-intensive operations such as sequencers [11], lock services [101, 56], and synchronization barriers [99], which often involve many-to-one contention on a small set of hot memory locations. Furthermore, most data-intensive applications use pointer-based data structures such as graphs, unbalanced trees, unstructured grids, and sparse matrices. These structures generate fine-grained and irregular memory access patterns [55]. As discussed in Section II-A, traditional DMA mechanisms are unsuitable for such access patterns. In contrast, CXL.cache efficiently supports such workloads through low-latency, coherent access to shared memory.
@@ V-B Remote Procedure Calls
Remote procedure calls (RPCs) serve as the fundamental inter-process communication layer for modern cloud services, supporting diverse distributed applications (e.g., microservices [62], cloud storage [26], and machine learning [53]). An RPC abstracts remote compute resources by making a function call to a remote machine similar to a local call. Its underlying stack (typically implemented as a user-space library [30, 73]) transparently handles connection management, network protocols, parameter (de)serialization, and thread scheduling. Industry measurements reveal significant RPC overheads: Google reports that RPCs account for 7.1% of CPU cycles in production clusters [69]; Meta observes that parameter (de)serialization alone consumes 6.7% of CPU cycles in seven critical microservices [76]. These findings highlight the importance of accelerating the RPC stack especially (de)serialization to reclaim CPU resources for application workloads.
@@ V-B1 Conventional RPC offloading design in PCIe-NIC
Among existing serialization libraries [63, 9, 23], we focus on Google’s Protobuf [63] in this paper, as it has widespread adoption in today’s cloud applications. Two key characteristics of RPC messages critically impact acceleration designs. (1) The vast majority of messages are extremely small; a Google cluster analysis shows that 56% of messages are bytes and 93% are bytes [39]. (2) Nested messages can be defined via pointer references, analogous to pointer chasing, incurring significant cumulative overhead during (de)serialization[62]; in real-world applications, the nesting of RPC message may exceed ten levels [39, 69].
A recent study RpcNIC [98] proposed a hardware offloading design of the RPC stack to a PCIe-NIC, which we have implemented in SimCXL as shown in Fig. 10. The host pre-runs the Protobuf compiler to store message structure metadata in a schema table, which guides message fields to decode in in-memory C++ objects or encode them into binary sequences. When an RPC request arrives, the NIC deserializer decodes it field-by-field, accumulating results in a on-chip temp buffer. When the current request’s deserialization is finished or the buffer is full, a one-shot DMA transfer to host memory is triggered. The NIC then increments the head pointer of a ring buffer in main memory through a DMA write to signal the arrival of a new request; the CPU processes the request and increments the tail pointer to indicate completion. For RPC responses, RpcNIC employs a pre-serialization mechanism; the CPU first invokes the on-chip memory copy engine (i.e., Intel’s DSA) to iteratively copy noncontiguous fields into a pre-allocated DMA-safe buffer. The CPU then updates an NIC-resident ring buffer via MMIO to signal the completion of pre-serialization. The NIC copies the prepared data from the host memory via a DMA read. The hardware serializer iteratively encodes the data in the NIC memory and transmits the RPC response over the network.
@@ VI-A1 Hardware testbed
Our hardware testbed is a high-performance dual-socket server, detailed in Table I. Each socket houses an Intel Gen4 Xeon Platinum 8468V processor, which integrates four CPU chiplets with each paired with a memory controller driving two memory channels. Each channel is populated with two DDR5 DIMMs. The four chiplets can be configured as a monolithic CPU or two/four NUMA nodes in sub-NUMA cluster mode (SNC) [1]. To ensure accurate and stable D2H memory access measurements, we enabled SNC-4 in the BIOS, dividing each socket into four separate NUMA nodes. Memory accesses from the same physical core to different NUMA nodes incur varying latencies depending on the number of NoC and UPI routing hops, thus manifesting the NUMA effect. The CXL devices on the testbed include an Intel Agilex I-Series FPGA Development Kit [2] and a Samsung memory expander [14]. The FPGA operates at and supports both CXL type-1/2 device configurations (denoted as CXL-FPGA) and standard PCIe configurations (denoted as PCIe-FPGA), connected to Socket1 via PCIe5.0 ×16. The CXL-FPGA integrates a hard CXL IP [15] and is equipped with a 4-way HMC, whereas the PCIe-FPGA integrates a multi-channel DMA IP [54]. The Samsung expander, connected to Socket1 via PCIe5.0 ×8, is exposed to the system as a CPU-less NUMA node.
@@ VI-A2 SimCXL simulator
SimCXL is developed based on gem5 v23.1, with X86O3CPU and of DDR5 dual-channel memory. In terms of CXL devices, the device driver is developed on top of the Linux v6.12 kernel. The modular CXL controller in SimCXL can model CXL type-1/2/3 devices. Device memory can directly leverage gem5’s native DDR, NVM, and HBM models; device compute units are customized following the gem5 programming framework. For type-1/2 devices (denoted as CXL-FPGAsim), a 4-way HMC is implemented in the Ruby subsystem. Based on the latency breakdowns presented in the CXL specification [18] and related studies [70, 36, 78, 81], the CXL controller and Ruby’s heterogeneous MESI protocol set multiple configurable parameters for tuning H2D and D2H access latency and bandwidth. For PCIe devices (denoted as PCIe-FPGAsim), we implemented DMA read and write engines to handle H2D and D2H data transfers, respectively. Both the CXL and PCIe device models can be configured to to match the FPGA configurations. Meanwhile, the frequency can also be set to to model real-world ASIC devices (denoted as CXL-ASICsim and PCIe-ASICsim, respectively), by frequency scaling based on the clock cycles measured from CXL-FPGA tests; this allows us to evaluate the realistic offloading performance of RAO and RPC on future production-grade CXL devices.
@@ VI-A4 Hardware calibration methodology
In this work, we calibrate the latency and bandwidth of CXL.cache for D2H memory accesses. The calibration of H2D memory access via CXL.mem will be presented in another paper. By tuning the model parameters of SimCXL, we are able to calibrate it to make its performance match that of CXL-FPGA. Note that the DMA performance calibration for SimCXL and PCIe-FPGA follows a similar approach. For D2H accesses via CXL.cache, cachelines may hit in the device’s HMC, host LLC, or host memory. HMC hits are tested by repeating address sequences. For LLC hits and memory hits, after initializing the target host memory addresses, we use the CLDEMOTE [12] instruction to push cachelines to the LLC or the CLFLUSH [13] instruction to flush cachelines to memory, followed by the LSU test sequence. For predictable performance, hyper-threading and hardware prefetchers are disabled, and CPU frequency is fixed at 2.4GHz during test.
@@ VI-C SimCXL Bandwidth Calibration
To measure the bandwidth of CXL.cache, we configured the LSU to issue a large number of memory requests until a stable bandwidth value was achieved. We observed that when the request count exceeds 512 (totaling ), the bandwidths for HMC hit, LLC hit, and memory hit all converge. We then report the bandwidth results in Fig. 15 at 2,048 requests (totaling ), which are averaged over 1,000 tests. On the 400 MHz CXL-FPGA, the peak bandwidths for HMC hit, LLC hit, and memory hit stabilize at , , and , respectively. Given the FPGA’s theoretical maximum bandwidth of at , LLC and memory hits reach only 55% and 52.7% of the peak bandwidth. This degradation probably stems from pipeline bubbles introduced by cache-coherence checks during requests routed to the host, a behavior also observed in CPU accesses to remote NUMA nodes [78, 81]. In contrast, HMC accesses, eliminating the need for host-side coherence checks, achieve 97.7% of the theoretical bandwidth. One can see that SimCXL accurately reproduces the bandwidth of CXL-FPGA at .
@@ VI-D Evaluation of Remote Atomic Operations
We evaluated six RAO access patterns from the CircusTent benchmark [88] on both PCIe-NIC and CXL-NIC in SimCXL, with throughput speedups shown in Fig. 17. The CENTRAL pattern, a many-to-one workload that models a distributed lock service, achieves the highest speedup (40.2) by caching the hotspot data in CXL-NIC’s HMC, avoiding costly PCIe DMA transfers. STRIDE1 follows with a 22.4 speedup, as its sequential atomic updates on elements leverage cacheline granularity, enabling cache hits for seven subsequent operations after fetching a line. SCATTER and GATHER patterns, which involve randomized updates in the global address space, exhibit moderate speedups due to lower cache hit rates. The SG workload, which combines SCATTER and GATHER, follows a similar trend. Even the fully random pattern RAND, with a near-zero cache hit rate, delivers a 5.5 throughput gain due to the lower access latency of CXL.cache compared to PCIe-based DMA for fine-grained accesses. These results demonstrate that CXL-NIC significantly enhances the distributed RAO throughput across various access patterns by leveraging low-latency memory access and on-device caching of hot data.
@@ VI-E Evaluation of Remote Procedure Calls
We evaluated de/serialization time for CXL-based RPC (CXL-NIC.*) versus PCIe-based RPC (RpcNIC [98]) using HyperProtoBench [39]. Fig. 18(a) shows deserialization results, with CXL-NIC achieving speedups of 1.33 (Bench5) to 2.05 (Bench1) in six benchmarks. This acceleration comes primarily from the CXL-NIC’s low-latency fine-grained writes. Although Bench5’s large string fields favor DMA, CXL.cache still outperforms RpcNIC by eliminating redundant copies and software overhead. Bench1, with small-field messages, achieves the highest speedup. Moreover, CXL-NIC uses NC-P instructions to push fields directly into the LLC, improving subsequent CPU access performance.
@@
| Simulator/Emulator | Cohet Support | CXL.cache Support | CXL.mem&io Support | CXL XPU Models | Full System | Hardware Calibration | Configurability | Sim. Error | Sim. Speed |
| CXLMemSim [92] | No | No | No | No | No | No | Medium | High | Medium |
| CXL-DMSim [86] | No | No | Yes | No | Yes | Yes | High | Low | Low |
| Mess+gem5 [21] | No | No | No | No | No | No | High | Medium | Low |
| QEMU [20] | No | No | Yes | No | Yes | No | High | High | High |
| Remote NUMA [74, 97] | No | No | No | No | No | N/A | Low | High | High |